OpenStudy (anonymous):

Find all real solutions of the quadratic equation. (If there is no real solution, enter write no real solution.) 7x2 + 14x − 1 = 0

OpenStudy (anonymous):

a = 7 b = 14 c = -1 *pLug them into the formula

OpenStudy (anonymous):

\[\frac{ 14\pm \sqrt{12.96} }{ 2 }\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

its -14

OpenStudy (anonymous):

not 14

OpenStudy (anonymous):

\[\frac{ -14 \pm \sqrt{(14)^2 - 4(7)(-1)} }{ 14 }\]

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

oo srry I made a mistake, I still got it wrong though. This is not correct

OpenStudy (anonymous):

\[\frac{ -14 \pm \sqrt{224} }{ 14 }\]

OpenStudy (anonymous):

dont u have to find the square root of 224 so it should be -14 pm square root of 14.96 over 14

OpenStudy (anonymous):

can you put in numbers what ur trying to say?

OpenStudy (anonymous):

\[ \frac{ -14\pm \sqrt{14.96} }{14}\]

OpenStudy (anonymous):

\[\frac{ -14 \pm 12\sqrt{61} }{ 14 }\]

OpenStudy (anonymous):

howd u get that?

OpenStudy (anonymous):

\[\sqrt{244} = 12\sqrt{61}\]

OpenStudy (anonymous):

Thank you for your help.

OpenStudy (anonymous):

no problem